# Quantum Wave Simulation with Sources and Loss Functions

### v1

@@ A.2 Proof of Theorem 2
The number of Pauli observables in equation (56) depends on $M$, as |$|\{I,X\}^{\otimes log_{2}M}|=M$. Therefore, the number of unitary observables that need to be estimated is $2M$. For verification, with $M=2$, $\mathbf{x}_{1}=\mathbf{a}$, and $\mathbf{x}_{1}=-\mathbf{b}$, we see that $\hat{\mathbf{O}}_{\Sigma}=\hat{\mathbf{O}}$ and $\hat{\mathbf{P}}_{\mathcal{S},\Sigma}=\hat{\mathbf{P}}_{\mathcal{S}}$, confirming that $l_{\Sigma,\mathcal{S}}^{2}=l_{2,\mathcal{S}}^{2}$.
Appendix B General Source Initialization
By definition, an isotropic radiator is a point source, which radiates the same intensity of radiation in all directions [73]. Examples of isotropic radiators include acoustic pressure sources or some radial vector field sources. In this section, we first show that an isotropic radiator acting in a homogeneous and isotropic domain generates a wave field, which satisfies rotational covariance around the point source location (Section B.3). We then utilize that rotationally covariant scalar and vector fields can be efficiently initialized on QCs (Lemma 4). Finally, we show that this initialization is also efficient if the homogeneous and isotropic volume around the point source is limited to radius $r_{s}$, as long as source activation time does not scale with the system size. This establishes a general implementation method for isotropic radiators, which does not impact the end-to-end runtime of the forward model, and which only demands material restrictions in a small volume around the source. The method extends beyond homogeneous materials and even beyond isotropic radiators as long as rotation covariance is satisfied.
B.1 Homogeneous and Isotropic Medium
For concreteness, here we choose the acoustic wave equation in $3D$. We consider a point source located at $\mathbf{x}_{0}$ within a homogeneous and isotropic domain. This source is active for a finite duration $T$ and is characterized by a source-time function $f(t)$. The 3D acoustic free-space homogeneous Green’s function reads [3]
$G_{3D,u}^{hom}(t,r)=\frac{1}{4\pi c^{2}r}\delta\left(t-t_{0}-\frac{r}{c}\right),\quad r=\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}+(z-z_{0})^{2}},$ (58)
where $\delta(t)$ denotes the Dirac delta function. Clearly, $G_{3D,u}^{hom}$ exhibits radial covariance, meaning that its scalar solution remains constant for any fixed radius $r$.
Next, let us consider the particle velocity solution of the Green’s function in the same setting, where the particle velocity is denoted by $\mathbf{v}$. The relationship between the particle velocity $\mathbf{v}$ and the pressure field $u$ in a source-free, homogeneous, and isotropic acoustic medium is given by
$\mathbf{v}(\mathbf{x},t)=-\frac{1}{\rho}\nabla\int_{-\infty}^{t}u(\mathbf{x},\tau)\,\mathrm{d}\tau,$ (59)
where $\rho$ is the density of the medium. Substituting the pressure Green’s function $G_{3D,u}^{hom}(t,r)$ into the above equation, we obtain
$\mathbf{v}(\mathbf{x},t)=-\frac{1}{\rho}\nabla\left(\frac{1}{4\pi c^{2}r}H\left(t-t_{0}-\frac{r}{c}\right)\right),$ (60)
where $H(t)$ is the Heaviside step function. Taking the spatial gradient, we have
$\nabla\left(\frac{1}{r}H\left(t-t_{0}-\frac{r}{c}\right)\right)$ $=\nabla\left(\frac{1}{r}\right)H\left(t-t_{0}-\frac{r}{c}\right)+\frac{1}{r}\nabla H\left(t-t_{0}-\frac{r}{c}\right)$ (61)
$=-\frac{\hat{\mathbf{r}}}{r^{2}}H\left(t-t_{0}-\frac{r}{c}\right)-\frac{\hat{\mathbf{r}}}{cr}\delta\left(t-t_{0}-\frac{r}{c}\right),$ (62)
where $\hat{\mathbf{r}}=\frac{\mathbf{r}}{r}$ is the unit radial vector. Substituting back, the particle velocity becomes
$\mathbf{v}(\mathbf{x},t)=-\frac{1}{\rho}\left(-\frac{\hat{\mathbf{r}}}{4\pi c^{2}r^{2}}H\left(t-t_{0}-\frac{r}{c}\right)-\frac{\hat{\mathbf{r}}}{4\pi c^{3}r}\delta\left(t-t_{0}-\frac{r}{c}\right)\right).$ (63)
Simplifying, the vectorial Green’s function for the particle velocity is
$\mathbf{G}_{3D,v}^{hom}(t,r)=\frac{\hat{\mathbf{r}}}{4\pi\rho c^{2}r^{2}}H\left(t-t_{0}-\frac{r}{c}\right)+\frac{\hat{\mathbf{r}}}{4\pi\rho c^{3}r}\delta\left(t-t_{0}-\frac{r}{c}\right).$ (64)
We see that the particle velocities also exhibit rotational covariance, with orientations radial to the point source at a fixed $r$. Consequently, the resulting time convolutions $\mathbf{K}_{u}=f*G_{3D,u}^{hom}$, which describes the initial pressure field, and $\mathbf{K}_{v}=f*\mathbf{G}_{3D,v}^{hom}$, which describes the initial particle velocity field, inherit the radial covariance and invariance properties, as established in Section B.3. As a result, we can initialize this wave field efficiently on a QC with Lemma 4. Note, that the particle velocity vectors have radial orientation and hence a single dimensional component of the vector contains all information needed for the construction of the full wave field. Hence, we obtain an initialization complexity of $\tilde{\mathcal{O}}(N^{1/D})$ quantum gates in the general $D$-dimensional case, which is independent of the vector dimension $C$.
B.2 Homogeneous Isotropic Medium in a Volume of Radius $r_{s}$
Finally, let us consider a setting, where the material around the point source is only homogeneous and isotropic up to a radius $r_{s}$ and where the source activation time $T$ is greater than the travel time of the wave in the homogeneous domain $t_{s}=r_{s}/c$. Here, the wave leaves the homogeneous domain before the source terminates. Hence, the Green’s function solution to this system is usually no longer analytically derivable and it does not satisfy rotational covariance around the source location. In this context, the previously introduced source implementation method is no longer applicable in its simple form. However, as long as $r_{s}$ is not vanishing, we can extend the method to this setting, as we will now show.
We will first discuss how the source can be decomposed, leading to multiple initial conditions. Then, we address the issue of evolving these initial conditions that start at different times, using a time-dilating Hamiltonian so that they all reach the same global time. Once all the initial conditions have reached the same time, they can be simulated in parallel under the same Hamiltonian with zero overhead toward the final time of simulation, $T$. The entire workflow is depicted in Figure 1. Finally, Theorem 2 allows us to sum the individual wave fields to compare their superposition to a target state or to estimate subspace energies.
Figure 1: Algorithm implementing a point source function on QCs. $A$ shows the source-time function for a single point location. We decompose the source-time function in time ($B$), which gives a set of functions that, in superposition, equate to the original source-time function ($C$). By convolving them in time with the homogeneous Green’s function ($D$), we obtain multiple sets of initial conditions at different global times ($E$). We evolve these initial conditions over different time intervals to the final simulation time $T$ ($F$). Here, the superposition of the source slices is equal to the solution of the wave equation ($G$). We can utilize this to efficiently estimate an $l_{2}$-distance to a measurement state using Theorem 2.
The point-source function $f(t)$ can be split into $J$ windows by multiplication with time window functions $W_{j}(t-\tau_{j})$, where $\tau_{j}$ is the time shift of the $j^{\text{th}}$ window. There exists multiple approaches of defining $W_{j}(t)$ such that
$f(t)\approx\sum_{j=1}^{J}W_{j}(t-\tau_{j})f(t),$ (65)
is a sufficiently accurate approximation. Defining the window functions and time shifts in a way such that the initial condition in both $u$ and $\mathbf{v}$ are smooth after multiplication is crucial for eliminating numerical dispersion effects in the subsequent HS. For the acoustic problem, this can be achieved with a symmetric double sigmoid function
where $z$ defines the steepness of the sigmoid functions. This function realizes overlapping windows that satisfy equation (65) with adaptable accuracy. In contrast to, e.g., box window functions ($z\rightarrow\infty$), sigmoid functions lead to smooth initial conditions. However, because the sigmoid functions do only converge to 00 at the window boundary in the limit where $z\rightarrow\infty$, choosing a finite $z$ provides a trade-off between source approximation accuracy and smoothness of the sections. All $J$ source terms for $f(t)$ (see equation (65)) can be recognized as individual sources that satisfy $T_{j}^{end}\leq T_{j}^{hom}$, meaning that for none of them, their wave leaves the homogeneous domain before it terminates.
We can now represent each $f(t)$ in terms of $J$ initial conditions as
$u_{j}(\mathbf{x},t-\tau_{j})=G_{3D,u}^{hom}(\mathbf{x},t)*(W_{j}(t)f(t+\tau_{j}))\quad|\mathbf{x}|<r\quad j=1,...,J.$ (67)
These initial conditions now individually satisfy rotational covariance around $\mathbf{x}_{0}$ and can hence individually be implemented efficiently. The introduced method of Section 4.2 can now be used to synchronize them to the same global time $T$ and to estimate loss functions that superpose all $J$ wave fields, which is efficient if $r_{s}$ is independent of $N$.
B.3 Covariance and Invariance Inheritance
Given the convolution $\mathbf{K}=f*\mathbf{G}$ defined by
$\mathbf{K}(\mathbf{x},t)=\int_{0}^{t}f(\tau)\mathbf{G}(\mathbf{x},t-\tau),\mathrm{d}\tau,$ (68)
we aim to show that $\mathbf{K}$ inherits the rotational covariance and invariance properties of $\mathbf{G}$.
Rotational Covariance: Let $\mathbf{S}=\mathbf{R}\oplus\mathbf{I}{D-C}\in\mathrm{SO}(D)$, where $\mathbf{R}\in\mathrm{SO}(C)$ acts on the first $C$ dimensions. Using the rotational covariance of $\mathbf{G}$ and the linearity of convolution, we have
$\mathbf{K}\big{(}\mathbf{S}\mathbf{x},t\big{)}=\int_{0}^{t}f(\tau)\mathbf{G}\big{(}\mathbf{S}\mathbf{x},t-\tau\big{)},\mathrm{d}\tau\ =\int_{0}^{t}f(\tau)\mathbf{R}\mathbf{G}(\mathbf{x},t-\tau),\mathrm{d}\tau\ =\mathbf{R}\int_{0}^{t}f(\tau)\mathbf{G}(\mathbf{x},t-\tau),\mathrm{d}\tau\ =\mathbf{R}\mathbf{K}(\mathbf{x},t).$ (69)
Thus, $\mathbf{K}$ is rotationally covariant.
Invariance: For any $\mathbf{y}\in\mathbb{R}^{D-C}$, using the invariance of $\mathbf{G}$,
$\mathbf{K}(\mathbf{x}+\mathbf{y},t)=\int_{0}^{t}f(\tau)\mathbf{G}(\mathbf{x}+\mathbf{y},t-\tau),\mathrm{d}\tau\ =\int_{0}^{t}f(\tau)\mathbf{G}(\mathbf{x},t-\tau)\mathrm{d}\tau\ =\mathbf{K}(\mathbf{x},t).$ (70)
Therefore, $\mathbf{K}$ is invariant in the remaining $D-C$ dimensions. Combining both properties, we conclude that $\mathbf{K}$ inherits the rotational covariance and invariance of $\mathbf{G}$.
B.4 Proof of Lemma 4
Our goal is to efficiently prepare the quantum state |$\ket{\psi}$ that encodes the vector field $\mathbf{K}$ on an isotropic radially symmetric grid centered at $\mathbf{x}_{0}$.
Exploiting Rotational Covariance
Since $\mathbf{K}$ is rotationally covariant around $\mathbf{x}_{0}$ in the first $C$ dimensions and invariant in the remaining dimensions; that is, for any rotation $\mathbf{S}\in\mathrm{SO}(D)$ that decomposes as $\mathbf{S}=\mathbf{R}\oplus\mathbf{I}_{D-C}$, where $\mathbf{R}\in\mathrm{SO}(C)$ and $\mathbf{I}_{D-C}$ is the identity in $\mathbb{R}^{D-C}$, we have
$\mathbf{K}\big{(}\mathbf{S}(\mathbf{x}-\mathbf{x}_{0})+\mathbf{x}_{0}\big{)}=\mathbf{R}\,\mathbf{K}(\mathbf{x}),\quad\forall\mathbf{x}\in\mathbb{R}^{D}.$ (71)
@@ B.4 Proof of Lemma 4
    $\mathbf{x}_{a,\boldsymbol{\theta}}=\mathbf{x}_{0}+r_{a}\,\hat{\mathbf{u}}_{\boldsymbol{\theta}},$ (72)
    where $\hat{\mathbf{u}}_{\boldsymbol{\theta}}$ is a unit vector determined by the angular indices $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{D-1})$.
@@
Select a reference direction, such as along the first coordinate axis, and compute $\mathbf{K}\big{(}\mathbf{x}_{a,0}\big{)}$ for $a=1,2,\dots,A$, where
@@
and $\mathbf{e}_{1}$ is the unit vector along the first coordinate axis. This requires $\mathcal{O}(A)=\mathcal{O}(N^{1/D})$ classical evaluations of $\mathbf{K}$.
@@
|$\ket{\phi}=\frac{1}{\sqrt{\mathcal{N}^{\prime}}}\sum_{c=1}^{C}\sum_{a=1}^{A}K_{c}\big{(}\mathbf{x}_{a,0}\big{)}\ket{c}\otimes\ket{a},$ (74)
where $\mathcal{N}^{\prime}=\sum_{c,a}\left|K_{c}\big{(}\mathbf{x}_{a,0}\big{)}\right|^{2}$. Since the number of terms is $CA$, this step requires $\tilde{\mathcal{O}}(CA)=\tilde{\mathcal{O}}(CN^{1/D})$ quantum gates [74, 75, 76, 77].
@@
    Binary Representation of Angular Indices: Each angular coordinate $\theta_{i}$ is discretized into $\Theta$ divisions and represented using $k=\log_{2}\Theta$ qubits. For $\Theta=2^{k}$, the total number of angular qubits is $(D-1)k$.
@@
    Rotation Decomposition: To rotate the initial vector $\mathbf{K}\big{(}\mathbf{x}_{a,0}\big{)}$ to all possible directions $\hat{\mathbf{u}}_{\boldsymbol{\theta}}$, we decompose the rotation $\mathbf{R}_{\boldsymbol{\theta}}\in\mathrm{SO}(C)$ into a sequence of $C-1$ rotations. Specifically, we perform rotations in the $(i,i+1)$-planes for $i=1,\dots,C-1$:
@@
    Sequence of Controlled Rotations: For each plane $(i,i+1)$, we apply controlled rotation gates $\mathrm{CR}_{i,i+1}(\theta_{i}^{(m)})$, where $m$ indexes the bits in the binary representation of the angular coordinate $\theta_{i}$. The rotation angles $\theta_{i}^{(m)}$ are defined as:
@@
    $\mathrm{CR}_{i,i+1}(\theta_{i}^{(m)})=\ket{0}\bra{0}_{\theta_{i,m}}\otimes I^{\otimes C}+\ket{1}\bra{1}_{\theta_{i,m}}\otimes R_{i,i+1}(\theta_{i}^{(m)}),$ (78)
@@
The total number of controlled rotation gates required is $\mathcal{O}\left(Ck\right)$, since for each of the $C-1$ planes, we apply $k$ controlled rotations corresponding to the bits of the angular indices. Since $k=\log\Theta=\log A$, the total number of controlled rotations is $\mathcal{O}\left(C\log A\right)=\mathcal{O}\left(\frac{C}{D}\log N\right)$. However, for a fixed spatial dimension $D$ the gate complexity of implementing these rotations [78] is dominated by the complexity of initializing |$\ket{\phi}$, which requires $\tilde{\mathcal{O}}(CA)=\tilde{\mathcal{O}}(CN^{1/D})$ quantum gates [74, 75, 76, 77].
@@
$U_{\mathrm{rot}}\ket{\Phi}$ $=\frac{1}{\sqrt{\mathcal{N}^{\prime}\Theta^{D-1}}}\sum_{c^{\prime}=1}^{C}\sum_{a=1}^{A}\sum_{\boldsymbol{\theta}}\left(\prod_{i=1}^{C-1}R_{i,i+1}(\theta_{i})\right)_{cc^{\prime}}K_{c^{\prime}}(\mathbf{x}_{a,0})\ket{c}\otimes\ket{a}\otimes\ket{\boldsymbol{\theta}}$
$=\frac{1}{\sqrt{\mathcal{N}^{\prime}\Theta^{D-1}}}\sum_{c=1}^{C}\sum_{a=1}^{A}\sum_{\boldsymbol{\theta}}K_{c}(\mathbf{x}_{a,\boldsymbol{\theta}})\ket{c}\otimes\ket{a}\otimes\ket{\boldsymbol{\theta}},$ (79)
where we used the fact that $K_{c}(\mathbf{x}_{a,\boldsymbol{\theta}})=\left(\mathbf{R}_{\boldsymbol{\theta}}\mathbf{K}(\mathbf{x}_{a,0})\right)_{c}$ due to the rotational covariance property. This is the desired state |$\ket{\psi}$.
@@
Applying $U_{\mathrm{rot}}$ requires implementing rotations with angles that decrease exponentially with the number of qubits used to represent the angular coordinates. In practice, this poses challenges due to the finite precision of quantum gates. Specifically, the minimal rotation angle scales inversely with the state size $A$, resulting in a gate precision requirement that scales linearly with $A$. However, the precision requirement remains independent of $D$. This independence allows for the implementation of a larger total number of grid points $N=A^{D}$ without necessitating additional precision.
@@
We provide numerical implementations for the 2D and the 3D case in the supplements to the publication. Figure 2 shows the quantum circuits for the 2D and 3D case that construct the radially symmetric vector field from a single ray. Figure 3 shows the final output vector fields in 3D, where each vector element in the initialization ray points either in the $x$-, $y$-, or $z$-direction.
Conclusion
By exploiting the rotational covariance of $\mathbf{K}$ and the symmetry of the isotropic grid, we have shown that the quantum state |$\ket{\psi}$ can be initialized with $\mathcal{O}(N^{1/D})$ classical evaluations of $\mathbf{K}$ and $\tilde{\mathcal{O}}(CN^{1/D})$ quantum gates, where the tilde hides polylogarithmic factors. This marks a polynomial improvement for a spatial dimension $D>1$ over the conventional initialization, which requires at least $\tilde{\mathcal{O}}(CN)$ quantum gates if the number of auxiliary qubits used is kept constant [74, 75, 76, 77].
(a) X-direction
(b) Y-direction
(c) Z-direction
Figure 2: Initialized radially symmetric quantum state. The subplots show vector fields constructed from an initial state in $[1,0,0]$ direction with purely $X$-, $Y$-, or $Z$- vector components. These components can also be combined.

### v3

@@ A.2 Proof of Theorem 2
The number of Pauli observables in Eq. (58) depends on $M$, as |$|\{I,X\}^{\otimes\log_{2}(M)}|=M$. Therefore, the number of unitary observables that need to be estimated is $2M$. Consequently, the overall query complexity of the algorithm is $\mathcal{O}(M\log(1/\delta)/\epsilon)$. For verification, with $M=2$, $\mathbf{w}_{1}=\mathbf{a}$, and $\mathbf{w}_{2}=-\mathbf{b}$, we see that $\tilde{\mathbf{O}}_{\Sigma}=\tilde{\mathbf{O}}$ and $\tilde{\mathbf{P}}_{\mathcal{S},\Sigma}=\tilde{\mathbf{P}}_{\mathcal{S}}$, confirming that $l_{\Sigma,\mathcal{S}}^{2}=l_{2,\mathcal{S}}^{2}$.
A.3 Weighted $l_{2}$-norms
Corollary 5 (Weighted $l_{2}$-norm).
$l_{v}^{2}=\sum_{j}^{V}v_{j}l_{\mathcal{S}_{j}}^{2},$ (60)
Applying this to the comparison between two wave fields we obtain
$l_{B}^{2}$ $=\sum_{j}^{\bar{B}}B^{-1/2}_{j}||\mathbf{P}_{\mathcal{S}_{j}}\mathbf{w}_{Q}-\mathbf{P}_{\mathcal{S}_{j}}\mathbf{w}_{Q}^{\text{target}}||^{2}$ (61)
$=\sum_{j}^{\bar{B}}||\mathbf{P}_{\mathcal{S}_{j}}\mathbf{w}-\mathbf{P}_{\mathcal{S}_{j}}\mathbf{w}^{\text{target}}||^{2},$
where $\bar{B}$ is the number of unique weights in the diagonal matrix $\mathbf{B}$, i.e. unique material parameters, and $S_{j}$ is the subspace only containing grid points with material parameters $B_{j}$. Here, we need to assume that $\mathbf{B}$ is diagonal in $\mathcal{S}$, which holds true for the acoustic wave equation or the isotropic Maxwell’s equation. While this, in principle, allows for comparing two wave fields in the standard basis, we expect prohibitive sampling costs if the respective subspaces are small due to a loss of estimation globality (see Section VI.1).
Appendix B Sources
B.1 Proof of Lemma 3
Rotational Covariance
Since $\mathbf{w}^{sym}$ is rotationally covariant around $\mathbf{x}_{0}=0$, we have
$\mathbf{R}(\mathbf{S})\mathbf{w}^{sym}(\mathbf{x})=\mathbf{w}^{sym}(\mathbf{S}\mathbf{x}),\quad\forall\mathbf{x},\mathbf{S},$ (62)
where $\mathbf{S}\in\mathrm{SO}(D)$ uniquely decomposes $\mathbf{R}(\mathbf{S})\in\mathrm{SO}(C)$.
@@ B.1 Proof of Lemma 3
    $\mathbf{x}_{a,\boldsymbol{\theta}}=\mathbf{x}_{0}+r_{a}\,\tilde{\mathbf{u}}_{\boldsymbol{\theta}},$ (63)
    where $\tilde{\mathbf{u}}_{\boldsymbol{\theta}}$ is a unit vector determined by the angular indices $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{D-1})$.
@@
Select a reference direction, such as along the first coordinate axis, and compute $\mathbf{w}^{sym}(\mathbf{x}_{a,0})$ for $a=1,2,\dots,A$, where
@@
and $\mathbf{e}_{1}$ is the unit vector along the first coordinate axis. This requires $\mathcal{O}(A)=\mathcal{O}(N^{1/D})$ classical evaluations of $\mathbf{w}^{sym}$.
@@
|$\ket{\phi}=\frac{1}{\sqrt{\mathcal{N}^{\prime}}}\sum_{c=1}^{C}\sum_{a=1}^{A}w^{sym}_{c}\big{(}\mathbf{x}_{a,0}\big{)}\ket{c}\otimes\ket{a},$ (65)
where $\mathcal{N}^{\prime}=\sum_{c,a}\left|w^{sym}_{c}\big{(}\mathbf{x}_{a,0}\big{)}\right|^{2}$. Since the number of terms is $CA$, this step requires $\tilde{\mathcal{O}}(CA)=\tilde{\mathcal{O}}(CN^{1/D})$ quantum gates [77, 78, 79, 80].
@@
    Binary Representation of Angular Indices: Each angular coordinate $\theta_{i}$ is discretized into $\Theta$ divisions and represented using $k=\log_{2}(\Theta)$ qubits. For $\Theta=2^{k}$, the total number of angular qubits is $(D-1)k$.
@@
    Rotation Decomposition: To rotate the initial vector $\mathbf{w}^{sym}(\mathbf{x}_{a,0})$ to all possible directions $\tilde{\mathbf{u}}_{\boldsymbol{\theta}}$, we decompose the rotation $\mathbf{R}_{\boldsymbol{\theta}}\in\mathrm{SO}(C)$ into a sequence of $C-1$ rotations. Specifically, we perform rotations in the $(i,i+1)$-planes for $i=1,\dots,C-1$:
@@
    Sequence of Controlled Rotations: For each plane $(i,i+1)$, we apply controlled rotation gates $\mathrm{CR}_{i,i+1}(\theta_{i}^{(m)})$ in the real plane, where $m$ indexes the bits in the binary representation of the angular coordinate $\theta_{i}$. The rotation angles $\theta_{i}^{(m)}$ are defined as:
@@
    $\mathrm{CR}_{i,i+1}(\theta_{i}^{(m)})$ =|$=\ket{0}\bra{0}_{\theta_{i,m}}\otimes I^{\otimes C}$ (69)
    +|$+\ket{1}\bra{1}_{\theta_{i,m}}\otimes R_{i,i+1}(\theta_{i}^{(m)}),$
@@
The total number of controlled rotation gates required is $\mathcal{O}\left(Ck\right)$, since for each of the $C-1$ planes, we apply $k$ controlled rotations corresponding to the bits of the angular indices. Since $k=\log(\Theta)=\log(A)$, the total number of controlled rotations is $\mathcal{O}\left(C\log(A)\right)=\mathcal{O}\left(\frac{C}{D}\log(N)\right)$. However, for fixed dimensions $C$ and $D$ the gate complexity of implementing these rotations [81] is dominated by the complexity of initializing |$\ket{\phi}$, which requires $\tilde{\mathcal{O}}(CA)=\tilde{\mathcal{O}}(CN^{1/D})$ quantum gates [77, 78, 79, 80].
@@
$U_{\mathrm{rot}}\ket{\Phi}=$ $\frac{1}{\sqrt{\mathcal{N}^{\prime}\Theta^{D-1}}}\sum_{c^{\prime}=1}^{C}\sum_{a=1}^{A}\sum_{\boldsymbol{\theta}}\left(\prod_{i=1}^{C-1}R_{i,i+1}(\theta_{i})\right)_{cc^{\prime}}$ (70)
$w^{sym}_{c^{\prime}}(\mathbf{x}_{a,0})\ket{c}\otimes\ket{a}\otimes\ket{\boldsymbol{\theta}}$
$=$ $\frac{1}{\sqrt{\mathcal{N}^{\prime}\Theta^{D-1}}}\sum_{c=1}^{C}\sum_{a=1}^{A}\sum_{\boldsymbol{\theta}}$
$w^{sym}_{c}(\mathbf{x}_{a,\boldsymbol{\theta}})\ket{c}\otimes\ket{a}\otimes\ket{\boldsymbol{\theta}},$
where we used the fact that $w^{sym}_{c}(\mathbf{x}_{a,\boldsymbol{\theta}})=\left(\mathbf{R}_{\boldsymbol{\theta}}\mathbf{w}^{sym}(\mathbf{x}_{a,0})\right)_{c}$ due to the rotational covariance property. This is the desired state, |$\ket{\psi}$, which proves Lemma 3.
@@
Applying $U_{\mathrm{rot}}$ requires implementing rotations with angles that decrease exponentially with the number of qubits used to represent the angular coordinates. In practice, this poses challenges due to the finite precision of quantum gates. Specifically, the minimal rotation angle scales inversely with the state size $A$, resulting in a gate precision requirement that scales linearly with $A$. However, the precision requirement remains independent of $C$ and $D$. This independence allows for the implementation of a larger total number of grid points $N=A^{D}$ without necessitating additional precision.
@@
We provide numerical implementations for the 2D and the 3D case in the supplements to the publication. Figure 2 shows the quantum circuits for the 2D and 3D case that construct the radially symmetric vector field from a single ray.
B.2 Windowing
While different choices of windows are possible, in our implementation we use a symmetric double sigmoid function
where $z$ defines the steepness of the sigmoid functions. This function realizes overlapping windows that satisfy Eq. (36) with adaptable accuracy. In contrast to, e.g., box window functions ($z\rightarrow\infty$), sigmoid functions lead to smooth initial conditions. However, because the sigmoid functions do only converge to 00 at the window boundary in the limit where $z\rightarrow\infty$, choosing a finite $z$ provides a trade-off between source approximation accuracy and smoothness of the sections.
Note, that if the windows are chosen to be adjacent block functions with no overlap ($z=\infty$), their appliance to a source time function $f(t)$ does not increase the number of nonzero values, which need to be initialized. In this case we would obtain the same initialization complexity as for a non-windowed state. However, as we reduce the steepness of the windows, they begin to overlap. This increases the amount of initialized values and is, in the worst case, a multiplicative complexity increase with the number of windows $J$.